### 19 Apr 2026 v1

# Sequential $\Upsilon(nS)$ suppression in high-multiplicity pppp collisions: multi-differential constraints on hadronic mechanisms

Abstract
The multiplicity-dependent suppression of $\Upsilon(n\mathrm{S})$ states measured by CMS in $pp$ at $\sqrt{s}=7$ TeV [1], and of $\psi(2\mathrm{S})/J/\psi$ measured by LHCb at $\sqrt{s}=13$ TeV [2], is subjected to four multi-differential tests: cone isolation, azimuthal sectors, transverse sphericity, and prompt vs. non-prompt. Cone and sphericity close a scissors constraint: the local reading of the Comover Interaction Model is in tension with the cone data, its global reading with the sphericity data. The non-prompt flatness forces the mechanism to act at early proper times. None of the considered hadronic or string-based frameworks — CIM local or global, PYTHIA 8 MPI [11], rope hadronisation [10], CGC [9], Trainor TCM [12] — naturally satisfies the four constraints simultaneously in its published form. The surviving class is an early, globally correlated medium consistent with partonic degrees of freedom, co-occurring with the ALICE strangeness enhancement [13], the long-range ridge [17], and below the threshold of the partonic baryon–meson $v_{2}$ [15], in a density window compatible with the Campanini & Ferri equation of state [18].
1 Introduction
The question at the heart of this paper can be stated in one sentence: is the suppression of the excited $\Upsilon$ states observed in high-multiplicity proton–proton collisions driven by ordinary hadronic final-state scattering, or is it telling us that the smallest collision system at the LHC already produces a transient, partonic-like medium? For decades, the progressive melting of quarkonium bound states has been read as the canonical thermometer of colour deconfinement in heavy-ion physics [4, 5]: the larger and more loosely bound a state, the lower the temperature at which Debye screening dissolves it. Applied to bottomonium, this picture predicts a clean hierarchy in which $\Upsilon(3\mathrm{S})$, the fragile partner with a binding energy of only $\sim 200$ MeV, melts first; $\Upsilon(2\mathrm{S})$ next; and $\Upsilon(1\mathrm{S})$, tightly bound by more than an order of magnitude, last. The same hierarchy has now been measured by CMS at $\sqrt{s}=7$ TeV [1] and by LHCb at $\sqrt{s}=13$ TeV [2, 3] in plain $pp$ collisions — a system where, by textbook reasoning, no collective medium should form at all. The fact of the hierarchy is therefore not in question; what is in question is its origin.
Three explanations compete for the role. The first and most conservative is the Comover Interaction Model, in which the quarkonium is dissociated by inelastic scattering off the soft hadrons produced in the same rapidity region [6, 7, 8]. The same framework has been extended beyond conventional states to the exotic $X(3872)$ in high-multiplicity $pp$ [24], where it plays a central role in constraining the hadronic molecule versus compact tetraquark interpretation. It is natural: the denser the event, the more scatterers; the more scatterers, the more dissociation; the larger the state, the larger the cross section. The second ascribes the effect to initial-state or string-topology phenomena that modify the production step itself — Colour Glass Condensate at low $x$ [9], or the enhanced effective string tension of rope hadronisation [10] when overlapping strings from many multi-parton interactions rearrange. The third invokes a genuine, short-lived partonic droplet, with colour screening acting on the pre-resonance $b\bar{b}$ pair before the $\Upsilon$ state has even had time to form. The three pictures are not just different in name: they make different predictions for observables beyond the inclusive multiplicity trend, and this is what allows the data to discriminate between them.
The key is that both CMS and LHCb did not stop at measuring the ratios against the global multiplicity. They also resolved the events by the local track density around the $\Upsilon$ (cone isolation), by the direction of the extra activity with respect to the $\Upsilon$ (azimuthal sectors), by the global shape of the event (transverse sphericity), and — crucially — by the production mechanism itself, separating quarkonia born at the primary vertex from those produced in the displaced decays of $b$ hadrons. Each of these cuts is a test with a definite prediction. Local hadronic scattering must depend on local hadron density. A mechanism driven only by the total multiplicity cannot depend on the shape of the event at fixed multiplicity. A mechanism acting at late proper times must affect displaced quarkonia just as it affects prompt ones. The surprise, and the content of this paper, is that when one imposes all the cuts simultaneously every hadronic or string-based candidate fails at least one of them, while a short-lived partonic medium satisfies them jointly.
What makes this conclusion more than an isolated quarkonium argument is the context in which it sits. The onset of $\Upsilon$ suppression in $pp$ lines up, in the same multiplicity range, with the strangeness enhancement seen by ALICE [13] — the oldest and cleanest QGP signal known from heavy-ion physics — and with the long-range two-particle ridge that developed into a unique feature of high-multiplicity $pp$ data at the LHC [17]. At still higher multiplicities, ALICE has recently reported the baryon–meson $v_{2}$ grouping [15], the mass-ordering switch that is the fingerprint of partonic, rather than hadronic, collectivity. Independently, the equation-of-state analysis of Campanini & Ferri [18] places a soft-to-hard transition band precisely where all these phenomena switch on. A Bjorken-type estimate of the energy density at the top of the accessible multiplicity range lands comfortably above the lattice-QCD deconfinement crossover [20, 21], yet modestly enough that the absence of jet quenching in the same collisions is not a contradiction but a natural consequence of the path-length and sensitivity scales involved. Seen together, these are a convergent set of observations that single out a specific class of mechanisms and close off every other.
2 Notation and observables
Throughout the paper we use the following symbols and conventions.
$N_{\mathrm{track}}$
charged-track multiplicity, CMS acceptance $|\eta|<2.4$, $p_{T}>0.4$ GeV/$c$. The relation $N_{\mathrm{track}}\simeq 2.83\,dN_{\mathrm{ch}}/d\eta$ simply reflects the integration of $dN_{\mathrm{ch}}/d\eta$ over the $|\eta|<2.4$ window after the reconstruction and $p_{T}$-selection efficiencies have been folded in; the explicit conversion table ($N_{\mathrm{track}}$ measured $\leftrightarrow$ $N_{\mathrm{track}}$ true $\leftrightarrow$ $dN_{\mathrm{ch}}/d\eta$ for $p_{T}>0$) is provided by CMS in Ref. [1].
$R_{21},R_{31}$
production ratios normalised to the ground state, $R_{n1}=Y(\Upsilon(n\mathrm{S}))/Y(\Upsilon(1\mathrm{S}))$, $n=2,3$, for prompt yields $Y$.
$R_{32}$
the differential ratio $R_{32}=Y(\Upsilon(3\mathrm{S}))/Y(\Upsilon(2\mathrm{S}))=R_{31}/R_{21}$, which cancels common-mode systematics (muon trigger, reconstruction, correlated feed-down) and isolates the sequential hierarchy.
$N^{\Delta R}_{\mathrm{track}}$
tracks with $p_{T}>0.4$ GeV/$c$ inside a cone $\Delta R<0.5$ around the $\Upsilon$ flight axis, where $\Delta R\equiv\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}$.
$N^{\Delta\phi}_{\mathrm{track}}$
sector multiplicity as a function of azimuthal angle $\Delta\phi$ between a track and the $\Upsilon$: forward $|\Delta\phi|<\pi/3$, transverse $\pi/3<|\Delta\phi|<2\pi/3$, backward $|\Delta\phi|>2\pi/3$.
$S_{T}$
transverse sphericity, $S_{T}=2\lambda_{2}/(\lambda_{1}+\lambda_{2})$, from the two eigenvalues of the transverse momentum tensor; $S_{T}\to 0$ jet-like, $S_{T}\to 1$ isotropic.
$p_{T}^{\Upsilon}$
transverse momentum of the reconstructed $\Upsilon$.
The paper is organised as follows. Sec. 3 presents the multiplicity dependence of $R_{21}$ and $R_{31}$ inclusive and at high $p_{T}$. Sec. 4 uses the cone-isolation test to falsify local hadronic mechanisms. Sec. 5 uses the azimuthal-sector data to remove any residual near-side scenario. Sec. 6 uses the transverse sphericity to falsify every mechanism that is a function of $N_{\mathrm{track}}$ alone. Sec. 7 introduces the LHCb non-prompt $\psi(2\mathrm{S})/J/\psi$ temporal constraint. Sec. 8 gathers the verdicts on the hadronic and string-based frameworks. Sec. 9 connects the onset to strangeness enhancement, to the ridge, and to the ALICE partonic-flow observation. Sec. 10 estimates the Bjorken energy density and connects it to the equation-of-state analysis of Campanini & Ferri. Sec. 11 explains why the absence of jet quenching in $pp$ is compatible with a partonic droplet.
3 Suppression in multiplicity: inclusive and high-$p_{T}$
The basic experimental fact is the monotonic decrease of both $R_{21}$ and $R_{31}$ with $N_{\mathrm{track}}$ in both event classes measured by CMS (Fig. 1).
Figure 1: CMS $pp$ data at $\sqrt{s}=7$ TeV (own plot from HEPData of Ref. [1]). Top row: $R_{21}=Y(\Upsilon(2\mathrm{S}))/Y(\Upsilon(1\mathrm{S}))$. Bottom row: $R_{31}=Y(\Upsilon(3\mathrm{S}))/Y(\Upsilon(1\mathrm{S}))$. Left column: inclusive $p_{T}^{\Upsilon}>0$, $|y|<1.93$. Right column: $p_{T}^{\Upsilon}>7$ GeV/$c$, $|y|<1.2$. In all four panels the ratio decreases monotonically with $N_{\mathrm{track}}$; the high-$p_{T}$ sample covers a wider multiplicity range and shows a cleaner monotonic trend. $\Upsilon(3\mathrm{S})$ is always more suppressed than $\Upsilon(2\mathrm{S})$ (sequential suppression).
Two features matter for what follows.
First, the suppression is monotonic in $N_{\mathrm{track}}$ for both $R_{21}$ and $R_{31}$, and the relative suppression of $R_{31}$ is stronger than that of $R_{21}$ — the expected sequential hierarchy by state size.
Second, at the same $N_{\mathrm{track}}$ the high-$p_{T}$ sample lies systematically above the inclusive one. Reading Fig. 1 at corresponding multiplicity bins, the high-$p_{T}$ ratios are $R_{21}(p_{T}>7)/R_{21}(p_{T}>0)\simeq 1.2$ and $R_{31}(p_{T}>7)/R_{31}(p_{T}>0)\simeq 1.4$ across the multiplicity range accessible to both samples. In other words, requiring a fast $\Upsilon$ reduces the amount of suppression, and reduces it more strongly for the more loosely bound state. This is the qualitative signature one expects from a kinematic escape: a fast quarkonium crosses the dense region in a time $\tau_{\mathrm{cross}}\sim R/(\beta\gamma)$ that shortens with $p_{T}^{\Upsilon}$, so a larger fraction of the population escapes intact, and the escape fraction is naturally higher for the more fragile states.
A fully differential picture of the effect is provided by the CMS $p_{T}^{\Upsilon}$-sliced measurement [1], which we replot in Fig. 2.
Figure 2: CMS $pp$ data at $\sqrt{s}=7$ TeV [1]: $R_{21}$ (left) and $R_{31}$ (right) as a function of $N_{\mathrm{track}}$ in six exclusive $p_{T}^{\Upsilon}$ slices, $0<p_{T}^{\Upsilon}<5$, $7<p_{T}^{\Upsilon}<9$, $9<p_{T}^{\Upsilon}<11$, $11<p_{T}^{\Upsilon}<15$, $15<p_{T}^{\Upsilon}<20$, $20<p_{T}^{\Upsilon}<50$ GeV/$c$ (from dark to light, all for $|y|<1.2$). Own plot from the public HEPData tables of Ref. [1]. At fixed $N_{\mathrm{track}}$ the slices are clearly ordered by $p_{T}^{\Upsilon}$: higher momentum gives systematically higher $R_{n1}$. The $N_{\mathrm{track}}$-dependence itself evolves with $p_{T}^{\Upsilon}$: the low-$p_{T}$ slices show the strongest decrease with multiplicity, while the $20<p_{T}^{\Upsilon}<50$ slice is compatible with flat behaviour. The effect is more pronounced for $R_{31}$ than for $R_{21}$.
Two observations follow directly from the figure: (i) at fixed $N_{\mathrm{track}}$ the six slices are ordered from bottom to top by increasing $p_{T}^{\Upsilon}$ for both $R_{21}$ and $R_{31}$, so the $p_{T}$-dependent survival is not an artefact of inclusive-sample averaging; (ii) the slope of $R_{n1}$ versus $N_{\mathrm{track}}$ weakens monotonically as $p_{T}^{\Upsilon}$ increases, to the point that the highest-$p_{T}$ slice is compatible with no $N_{\mathrm{track}}$-dependence within uncertainty, as one expects if the $\Upsilon$ crossing time becomes much shorter than the dissociation timescale. This is not a measurement of a specific dissociation cross section, but it is the differential evidence, in the CMS data themselves, for a $p_{T}^{\Upsilon}$-dependent survival mechanism. The LHCb $\psi(2\mathrm{S})/J/\psi$ analysis of Sec. 7 provides a consistent $p_{T}$-resolved view on charmonium.
Third, the mean transverse momentum of the three $\Upsilon$ states as a function of $N_{\mathrm{track}}$ carries the same message in a complementary, inclusive-yield-weighted form (Fig. 3).
Figure 3: Mean transverse momentum $\langle p_{T}^{\Upsilon}\rangle$ of the three $\Upsilon(n\mathrm{S})$ states as a function of $N_{\mathrm{track}}$, separately for the inclusive sample ($p_{T}^{\Upsilon}>0$, $|y|<1.93$, left) and the high-$p_{T}$ sample ($p_{T}^{\Upsilon}>7$ GeV/$c$, $|y|<1.2$, right). Own plot from the public HEPData tables of Ref. [1]. At low $N_{\mathrm{track}}$ the three means coincide, confirming that the three states share the same production kinematics. As $N_{\mathrm{track}}$ increases, the means split into the hierarchy $\langle p_{T}\rangle(\Upsilon(3\mathrm{S}))>\langle p_{T}\rangle(\Upsilon(2\mathrm{S}))>\langle p_{T}\rangle(\Upsilon(1\mathrm{S}))$. The same ordering is present in both kinematic samples; the absolute scale is set by the $p_{T}$ cut.
At the lowest accessible multiplicity the three distributions have essentially the same mean $p_{T}^{\Upsilon}$: the three states are produced with the same kinematics. As $N_{\mathrm{track}}$ grows, the three means fan out and order themselves into $\langle p_{T}\rangle(\Upsilon(3\mathrm{S}))>\langle p_{T}\rangle(\Upsilon(2\mathrm{S}))>\langle p_{T}\rangle(\Upsilon(1\mathrm{S}))$. The ordering by state size is the same in the inclusive and in the high-$p_{T}$ sample; only the absolute scale changes with the kinematic cut. This is consistent with a picture in which a dissociative mechanism depletes each state preferentially at low $p_{T}^{\Upsilon}$, so the surviving $\Upsilon$ population gets progressively “pushed” to higher $p_{T}$; and since the more loosely bound states are depleted more strongly, their surviving mean shifts upward more. We stress that Fig. 3 is a shape statement about the surviving population — not an independent dynamical measurement — and that alternative production-level explanations for the same ordering cannot be ruled out on inclusive yields alone. Taken together with Fig. 1, however, the two independent readings of the data point in the same direction: a $p_{T}^{\Upsilon}$-dependent survival, stronger for the more loosely bound states. These two features set the stage but do not discriminate between the candidate mechanisms. The discrimination comes from the differential tests of the following sections.
4 Cone isolation constrains local hadronic CIM
The test.
CMS [1] splits events at fixed $N_{\mathrm{track}}$ into two extreme classes according to the number of tracks inside a cone $\Delta R<0.5$ around the $\Upsilon$:
  • empty cone, $N^{\Delta R}_{\mathrm{track}}=0$, no charged track (above $p_{T}>0.4$ GeV/$c$) collinear with the $\Upsilon$;
  • dense cone, $N^{\Delta R}_{\mathrm{track}}>2$, at least three such tracks.
Any mechanism in which the $\Upsilon$ is dissociated by inelastic scattering with nearby hadrons must produce different suppression in the two classes: more local scatterers $\Rightarrow$ stronger suppression. The sign of the expected effect is unambiguous.
The data.
Figure 4: Cone-isolation test at $\sqrt{s}=7$ TeV. Left: $R_{21}$. Right: $R_{31}$. Blue circles: empty cone ($N^{\Delta R}_{\mathrm{track}}=0$). Red diamonds: dense cone ($N^{\Delta R}_{\mathrm{track}}>2$). Own plot from HEPData of Ref. [1], $p_{T}^{\Upsilon}>7$ GeV/$c$, $|y|<1.2$. The two classes are statistically indistinguishable across $10\lesssim N_{\mathrm{track}}\lesssim 80$; any local hadronic mechanism predicts a clear separation that is not seen in the data.
At $N_{\mathrm{track}}\simeq 70$ the empty-cone suppression is $20.7\%$ relative to the lowest-$N_{\mathrm{track}}$ bin, the dense-cone value is $17.2\%$, and the two are compatible within uncertainty.
Consequence for hadronic CIM (local reading).
In the Armesto–Gavin–Ferreiro comover model with Cornell-potential radii $r_{1S}=0.226$ fm, $r_{2S}=0.486$ fm, $r_{3S}=0.690$ fm, the geometric dissociation cross sections scale as $\sigma_{2S}/\sigma_{1S}\simeq 4.6$, $\sigma_{3S}/\sigma_{1S}\simeq 9.3$, and the survival probability is $P_{\mathrm{surv}}(nS)=\exp[-\sigma_{nS}\,\rho_{\mathrm{com}}\,L]$. If $\rho_{\mathrm{com}}$ is identified with the local hadron density around the $\Upsilon$ — the physically honest reading of the word “comover” — the dense cone must produce substantially stronger suppression than the empty cone.
Keeping the three levels separate: data: the empty-cone and dense-cone classes are statistically indistinguishable within uncertainty. Experimental constraint: whatever mechanism generates the suppression does not scale with the local track density around the $\Upsilon$; a local-density proxy is not the right variable. Interpretation: in its local-density form the CIM is in direct tension with the cone data; the framework can of course be preserved by promoting $\rho_{\mathrm{com}}$ to a global quantity (Sec. 6 will show that this move is itself in tension with the sphericity-resolved data), but its most natural, physically transparent reading does not survive the cone test. This closes the first door.
5 Azimuthal independence removes the “near-side” escape
The cone test removes local scattering in a forward cone around the $\Upsilon$. A local mechanism could, however, be imagined where tracks collinear with the $\Upsilon$ drive the suppression while tracks in other directions do not. CMS resolves this by measuring the suppression as a function of the multiplicity restricted to three non-overlapping azimuthal sectors (forward, transverse, backward) relative to the $\Upsilon$ flight direction.
Figure 5: Suppression ratios as a function of sector multiplicity $N^{\Delta\phi}_{\mathrm{track}}$ in three azimuthal bins relative to the $\Upsilon$: forward ($|\Delta\phi|<\pi/3$), transverse ($\pi/3<|\Delta\phi|<2\pi/3$), backward ($|\Delta\phi|>2\pi/3$). Own plot from HEPData of Ref. [1], $p_{T}^{\Upsilon}>7$ GeV/$c$, $|y|<1.2$. Left: $R_{21}$. Right: $R_{31}$. All three sector datasets are statistically consistent within uncertainty: the suppression scales with the sector multiplicity at the same rate in every direction. Tracks on the opposite side of the $\Upsilon$ (backward) drive the suppression as efficiently as tracks collinear with it (forward). No near-side hierarchy survives this test.
The backward result is particularly instructive: backward tracks are geometrically far from the $\Upsilon$, so they cannot represent a local scattering environment. The fact that they correlate with the suppression at the same rate as forward tracks implies that what drives the suppression is a global activity variable, of which all three sector multiplicities are proxies. Azimuthal independence is therefore a necessary but not sufficient element: it rules out proximity-based mechanisms, but does not by itself rule out a mechanism in which $N_{\mathrm{track}}$ itself is the relevant variable. The next step discriminates this remaining possibility.
6 Sphericity constrains every $N_{\text{track}}$-only mechanism
The test.
At fixed $N_{\mathrm{track}}$, the transverse sphericity $S_{T}$ separates events with the same global multiplicity into radically different topologies:
  • $S_{T}<0.55$: jet-like events, dominated by a hard process with a strong back-to-back axis;
  • $S_{T}>0.85$: isotropic events, dominated by many soft MPI distributed over azimuth.
Any mechanism that parametrises the dissociation probability as a function of $N_{\mathrm{track}}$ alone must give the same suppression in the two classes. This includes: the hadronic global CIM ($\rho_{\mathrm{com}}\propto N_{\mathrm{track}}$), the Colour Glass Condensate (initial-state $Q_{s}^{2}$ set by the overall collision density), and pure PYTHIA MPI (no interaction with the produced $\Upsilon$).
The data.
Figure 6: Sphericity test: extreme classes. Left: $R_{21}$. Right: $R_{31}$. Blue circles: jet-like $S_{T}\in[0.00,0.55]$. Red diamonds: isotropic $S_{T}\in[0.85,1.00]$. Own plot from HEPData of Ref. [1], $p_{T}^{\Upsilon}>7$ GeV/$c$, $|y|<1.2$. In jet-like events both $R_{21}$ and $R_{31}$ are essentially flat in $N_{\mathrm{track}}$. In isotropic events both ratios decrease monotonically and strongly. The two populations are selected to have the same $N_{\mathrm{track}}$ but different topology; the fact that the suppression depends on topology and not just on $N_{\mathrm{track}}$ falsifies every $N_{\mathrm{track}}$-only mechanism.
Globally this corresponds to the two significances that can be read directly off the extreme classes shown in Fig. 6: the $R_{32}$-versus-$N_{\mathrm{track}}$ slope is compatible with zero in jet-like events ($S_{T}<0.55$, $1.9\sigma$) and reaches $5.0\sigma$ in isotropic events ($S_{T}>0.85$[1]. The sphericity filter therefore operates on both $R_{21}$ and $R_{31}$ and acts coherently on the differential ratio $R_{32}$.
Consequence.
Every mechanism parametrised by $N_{\mathrm{track}}$ alone — hadronic global CIM, CGC in its standard-form $Q_{s}^{2}(N_{\mathrm{track}})$ reading, pure MPI — is in direct tension with the sphericity-resolved data at fixed $N_{\mathrm{track}}$. Together with Sec. 4 this produces a symmetric closure: the local reading of the CIM is in tension with the cone test; the global reading with the sphericity test. No continuous parameter interpolates between the two without re-opening one of the failures. This is the scissors constraint: no simple reinterpretation of the comover density can simultaneously account for cone-independence and sphericity-dependence. We frame this as a joint constraint on the naturally expected forms of the CIM rather than as an exclusion of the model class as a whole: a comover-like mechanism that acquires an explicit dependence on event topology, or that couples to partonic rather than hadronic degrees of freedom, is not ruled out by this argument alone and could in principle survive the constraint.
On the differential ratio $R_{32}$ it is important to keep the three levels separated. Data: at $N_{\mathrm{track}}\simeq 74$ the high-sphericity class has $R_{32}\simeq 0.86\times R_{32}(\mathrm{jet})$, a $14\%$ gap between topologies at the same $N_{\mathrm{track}}$ [1]. Experimental constraint: the sequential hierarchy is driven by the event topology at fixed multiplicity rather than being inherited from production kinematics, because production kinematics are fixed when $N_{\mathrm{track}}$ is fixed. Interpretation, with caveats: the $R_{32}$ evidence is most natural in a core–corona picture where the isotropic core preferentially dissolves the more loosely bound state, but a mixing between core and corona contributions, or a non-trivial differential feed-down pattern, could in principle shift the quantitative picture. We do not rely on a specific CIM prediction for $R_{32}$ here: the observed topology-resolved $14\%$ gap is a direct experimental fact, and its interpretation in terms of any specific model requires inputs (Cornell radii, $\rho_{\mathrm{com}}(N_{\mathrm{track}},S_{T})$, feed-down fractions) that are not uniquely fixed.
7 Temporal constraint: LHCb non-prompt $\psi(2\mathrm{S})/J/\psi$
[…]
Figure 7: Normalised $\sigma_{\psi(2\mathrm{S})}/\sigma_{J/\psi}$ as a function of the non-dimensionalised charged-track multiplicity, $N^{\mathrm{PV}}_{\mathrm{tracks}}/\langle N^{\mathrm{PV}}_{\mathrm{tracks}}\rangle_{\mathrm{NB}}$, separately for prompt (orange) and non-prompt (blue) production. Values computed from the numerical Tables 4–8 of Appendix A of Ref. [2] (arithmetic mean of the double-differential ratios over the 15 cells in $(p_{T},y)$ of each multiplicity bin, then self-normalised to the mean across the five multiplicity bins). The co-mover model prediction [2] is shown as the hatched band. Error bars are approximate (statistical propagation of the per-cell uncertainties assuming uncorrelated errors); systematic uncertainties are not shown. The prompt ratio decreases monotonically and is well described by the co-mover model band except in the low-multiplicity region. The non-prompt ratio is flat within uncertainties across the full range.
[…]
A complementary reading is obtained by resolving the prompt and non-prompt ratios in slices of the charmonium transverse momentum $p_{T}^{\psi}$.
Figure 8: LHCb $pp$ data at $\sqrt{s}=13$ TeV: normalised $\psi(2\mathrm{S})/J/\psi$ cross-section ratio as a function of the non-dimensionalised multiplicity, in five exclusive $p_{T}^{\psi}$ slices ($0.3<p_{T}<2$, $2<p_{T}<4$, $4<p_{T}<6$, $6<p_{T}<8$, $8<p_{T}<20$ GeV/$c$), separately for prompt (left) and non-prompt (right) production. Own plot from the numerical Tables 4–8 of Appendix A of Ref. [2]; the self-normalisation is performed slice by slice to the average of the five multiplicity bins of the same slice. For prompt production the low-$p_{T}$ slices show the strongest multiplicity dependence, and the high-$p_{T}$ slice ($8<p_{T}<20$, yellow crosses) becomes compatible with flat within uncertainty. For non-prompt production no $p_{T}$ slice shows a significant multiplicity trend. This is the $\psi(2\mathrm{S})/J/\psi$ analogue of the CMS $\Upsilon$ evidence in Fig. 2.
Two observations parallel the CMS picture of Sec. 3: (i) among prompt charmonia, the suppression weakens as $p_{T}^{\psi}$ grows, so that the highest-$p_{T}$ slice is compatible with no multiplicity dependence within uncertainty — the kinematic-escape pattern already observed for $\Upsilon(n\mathrm{S})$; (ii) for non-prompt charmonia the ratio is flat irrespective of $p_{T}^{\psi}$, so the flatness reported in Fig. 7 is not the effect of integrating over $p_{T}$ but holds slice by slice. Together, Figs. 7 and 8 show that the $p_{T}^{\psi}$-dependence and the prompt/non-prompt dichotomy are two independent features of the LHCb data, each carrying a separate physical constraint: a finite-size medium (from the $p_{T}^{\psi}$ ordering) acting at early proper times (from the prompt-only trend).
One caveat is that the non-prompt argument is measured in charmonium, while the sequential suppression pattern of Sections 3 to 6 concerns bottomonium. The logical content we use here is not a flavour-specific statement but a temporal one: whatever mechanism produces the prompt-only multiplicity dependence must act at proper times short compared to the displacement of $b$-hadron decays, i.e. at the pre-resonance stage. This kind of temporal bound is model-independent and applies equally to any quarkonium, charmonium or bottomonium, since both are formed within $\mathcal{O}(1)$ fm/$c$ of the primary vertex. We therefore take the non-prompt flatness as a general constraint on the kind of mechanism at work rather than as direct evidence on $\Upsilon(n\mathrm{S})$ specifically.
8 Review of hadronic and string-based frameworks
We confront, one by one, the frameworks that have been proposed to explain the suppression in $pp$ with the four independent filters (cone, azimuth, sphericity, $p_{T}$-escape) plus the non-prompt flatness.
Local hadronic CIM.
Dissociation by nearby hadrons; cross sections scale as $(r_{nS}/r_{1S})^{2}$, so the model predicts more suppression in the dense-cone class than in the empty-cone class. The CMS data show the two classes to be indistinguishable within uncertainty. Verdict: in direct tension with the cone test; the model in its local-density reading does not reproduce the sign of the predicted effect.
Global hadronic CIM ($\rho_{\mathrm{com}}\propto N_{\mathrm{track}}$).
Prediction: cone- and azimuth-independent (satisfied), but shape-independent at fixed $N_{\mathrm{track}}$. Data: $R_{21}$ and $R_{31}$ both strongly shape-dependent (Fig. 6). Verdict: incompatible with the sphericity-resolved data in its $\rho_{\mathrm{com}}\propto N_{\mathrm{track}}$ form.
Pure MPI (PYTHIA 8, no extra physics) [11].
Multi-parton interactions generate $N_{\mathrm{track}}$ but do not interact with the produced $\Upsilon$. Prediction: $R_{n1}$ flat in $N_{\mathrm{track}}$. Data: monotonic decrease, enhanced in isotropic events. Verdict: outside the model’s predictive scope; PYTHIA 8 with only MPI is not designed to reproduce a multiplicity-dependent $\Upsilon(n\mathrm{S})$ hierarchy.
Colour reconnection / rope hadronisation [10].
Overlapping strings from many MPI form ropes with enhanced effective tension $\kappa_{\mathrm{eff}}>\kappa_{0}$, preferentially suppressing loosely-bound states. Cone: OK (string overlaps set at production time, not driven by which cone the fragments end up in). Azimuth: OK (strings approximately uniform in $\phi$). Sphericity: the direction of the effect is correct (isotropic events have more MPI and hence larger rope tension), but no published PYTHIA+rope simulation reproduces, at fixed parameters, the full jet-like (compatible with zero, $1.9\sigma$) vs. isotropic ($5.0\sigma$) gap observed by CMS. $p_{T}$-escape: the rope tension does not depend on $p_{T}^{\Upsilon}$, so the model has no native escape mechanism; the LHCb low-$p_{T}$ enhancement of suppression is not naturally reproduced. Verdict: qualitatively consistent with cone and azimuth, quantitatively unvalidated on sphericity, and structurally missing a native $p_{T}$-escape; not ruled out as a partial contribution but not a complete explanation on present evidence.
Colour Glass Condensate (initial-state saturation) [9].
Quarkonium dissociated by a high-density gluon field with saturation scale $Q_{s}^{2}$, set before the collision. Cone: OK. Azimuth: OK. Sphericity: a CGC calibrated by the overall initial-state density is $N_{\mathrm{track}}$-driven, and therefore predicts identical suppression in jet-like and isotropic events at fixed $N_{\mathrm{track}}$, contrary to the data. $p_{T}$-escape: no finite-size geometry, no native escape mechanism. Verdict: in tension with the sphericity-resolved data and structurally missing a $p_{T}$-escape structure in its standard form.
Two-component underlying-event parametrisations [12].
A recurring alternative reading of the high-multiplicity $pp$ data is Trainor’s two-component model (TCM), in which every spectrum is decomposed into a “soft” component (scaling with the number of participant partons) and a “hard” component (scaling with the number of hard MPI), without invoking a medium. Applied to quarkonium, TCM re-expresses $R_{n1}(N_{\mathrm{track}})$ as an interpolation between a soft plateau and a hard-dominated tail: a suppression emerges from the fact that the soft component carries a different effective $R_{n1}$ than the hard component, and their relative weight shifts with $N_{\mathrm{track}}$. Tuned per observable, the model can be made to track either $R_{21}$ or $R_{31}$ individually. The problem is that the differential structure of the CMS data fixes the $N_{\mathrm{track}}$-dependence of the weights in a way that is not self-consistent across cuts: at fixed $N_{\mathrm{track}}$, jet-like and isotropic events have very different soft/hard decompositions (jet-like: hard-component dominated; isotropic: soft-component dominated) so the TCM prediction is different for the two sphericity classes, but the magnitude of the split does not reproduce the sphericity-resolved data; likewise, cone-empty and cone-dense events cannot be selected to simultaneously match the observed near-equality in the cone test and the observed strong gap in the sphericity test. In practice no public TCM fit has reproduced the full multi-differential CMS set at fixed parameters. Verdict: no demonstrated predictive content once the full multi-differential set is imposed; the model retains flexibility but has yet to match the combined constraints.
Table 1: Verdict matrix. OK = the framework accounts for the feature naturally; Tension = structural difficulty in the framework’s standard/published form; Tunable = can be accommodated with parameter adjustments. No entry is meant as a statement that every variant of a framework is excluded; it reflects the status of the considered published formulations.
Framework Seq. suppr. Cone Sphericity $p_{T}$-escape Non-prompt flat
Local hadronic CIM Tension Tension Tension Tunable OK
Global hadronic CIM Tension OK Tension Tunable OK
Pure MPI (PYTHIA) Tension OK Tension Tension OK
CR / Rope hadronisation Tunable OK Tension Tunable OK
CGC (initial state) Tunable OK Tension Tension OK
Two-component UE Tension Tension Tension Tension OK
Globally correlated medium OK OK OK OK OK
None of the considered frameworks naturally and simultaneously accounts for the combined set of constraints in its published form. The entry in the last row is not a specific model but a description of the minimal structural features a surviving framework must have: it must act early (non-prompt flat), globally (sphericity), non-locally around the $\Upsilon$ (cone), and must carry a $p_{T}^{\Upsilon}$-dependent escape probability. The natural physical realisation of all four features simultaneously is an early, globally correlated medium with partonic degrees of freedom.
9 Multi-observable context: strangeness, ridge, partonic flow
The onset of $\Upsilon$ suppression in the CMS data is at $N_{\mathrm{track}}\approx 10$1515, corresponding to $dN_{\mathrm{ch}}/d\eta\approx 4$66 (using the conversion $N_{\mathrm{track}}\simeq 2.83\,dN_{\mathrm{ch}}/d\eta$). Three independent collective-like phenomena in the same collision system share the same onset window.
(a) Strangeness enhancement.
[…]
(b) Ridge: long-range near-side correlations.
Two-particle correlations in $\Delta\eta$$\Delta\phi$ in high-multiplicity $pp$ at the LHC show a near-side ridge at $\Delta\phi\simeq 0$ extending over several units of $\Delta\eta$ [17]. The ridge is a hallmark of early-time, long-range correlations in the initial-state gluon field, which persist through the evolution. The long-$\Delta\eta$ structure of the ridge is causally compatible only with an interaction at very short times, matching the non-prompt constraint of Sec. 7.
(c) Partonic flow (baryon–meson $v_{2}$ grouping).
ALICE 2024 [15] reports the first observation, with $\sim 5\sigma$ significance, of baryon–meson $v_{2}$ grouping in high-multiplicity $pp$ at $\sqrt{s}=13$ TeV. This is the mass ordering at low $p_{T}$ giving way to a baryon–meson ordering at intermediate $p_{T}$: the canonical signature of partonic collective flow with number-of-constituent-quark scaling. It requires the top $0.07\%$ of events in V0M multiplicity, $dN_{\mathrm{ch}}/d\eta\approx 26$ at midrapidity, and is absent in the low-multiplicity class.
Hierarchy of onsets.
Observable Onset ($dN_{\mathrm{ch}}/d\eta$)
$\Upsilon(2$S,3,3S)) sequential suppression $\approx 4$66
Strangeness enhancement $\approx 4$88
Baryon–meson $v_{2}$ grouping (partonic flow) $\gtrsim 17$2626
The $\Upsilon(n\mathrm{S})$ suppression begins before the explicit partonic-flow signal becomes measurable. This is not paradoxical: the loosely bound excited states have binding energies of a few hundred MeV ($\sim 200$ MeV for $\Upsilon(3\mathrm{S})$) and therefore are sensitive to a density regime in which partonic flow, set by the number of constituents participating coherently, is still below the detection threshold. The quarkonium is the earliest probe of the transition; partonic $v_{2}$ requires more developed collectivity before becoming measurable.
10 Bjorken energy density: consistency check
As a consistency check, not as a pillar of the argument, we estimate the Bjorken energy density [19] reached at the top of the multiplicity range and compare it to the lattice-QCD crossover band. We stress upfront that in $pp$ every input of this estimate is model dependent: the formation time $\tau_{0}$, the transverse area $A_{T}$, and the midrapidity $dE_{T}/dy$ are all inferred indirectly. The numbers below should be read as an order-of-magnitude benchmark, not as a measurement.
$\varepsilon_{\mathrm{BJ}}=\frac{1}{\tau_{0}A_{T}}\,\frac{dE_{T}}{dy}\Big|_{y=0},\qquad\tau_{0}\approx 1~\mathrm{fm}/c,~A_{T}=\pi R_{\mathrm{HBT}}^{2}\approx 10.2~\mathrm{fm}^{2},$ (1)
with $R_{\mathrm{HBT}}\approx 1.8$ fm measured in high-multiplicity $pp$ [1] and
$\frac{dE_{T}}{dy}\Big|_{y=0}\approx\langle m_{T}\rangle\times\tfrac{3}{2}\times dN_{\mathrm{ch}}/d\eta\approx 0.45\,\mathrm{GeV}\times 1.5\times 26\approx 17.6~\mathrm{GeV},$ (2)
where $\langle m_{T}\rangle\approx 0.45$ GeV is the transverse-mass average over the $\pi$, K, p spectra measured by ALICE at $\sqrt{s}=7$ TeV [16] and the factor $3/2$ accounts for neutral particles in the isospin-symmetric limit. Substituting:
$\varepsilon_{\mathrm{BJ}}\approx 1.7~\mathrm{GeV/fm}^{3},\qquad T\approx\left(\frac{30\,\varepsilon_{\mathrm{BJ}}}{37\pi^{2}}\right)^{1/4}\approx 180~\mathrm{MeV},$ (3)
using the massless-QGP Stefan–Boltzmann relation. The translation from $\varepsilon_{\mathrm{BJ}}$ to $T$ is itself crude because it assumes thermalisation and an ideal-gas equation of state, neither of which is established in $pp$.
Consistency with lattice QCD and with Campanini & Ferri.
This estimate exceeds the lattice-QCD pseudo-critical density $\varepsilon_{\mathrm{pc}}\approx 0.5$ GeV/fm3 [20, 21] by a factor $\sim 3$, with a temperature of the same order as $T_{\mathrm{pc}}\approx 156$ MeV. On independent grounds, the equation-of-state analysis of Campanini & Ferri [18], performed on $pp$ and $p\bar{p}$ multiplicity-dependent observables, identifies a soft-to-hard transition band $dN_{\mathrm{ch}}/d\eta\approx 6$2424. This band overlaps with the density window in which the signals discussed above become active. We read the convergence not as proof of deconfinement but as a consistency check: at the multiplicities where $\Upsilon(n\mathrm{S})$ sequential suppression, strangeness enhancement and (at higher $dN_{\mathrm{ch}}/d\eta$) the baryon–meson $v_{2}$ grouping all switch on, the naive hadronic description reaches its plausible limit of applicability [22]. The central claims of the paper rest on the multi-differential CMS and LHCb data analysed in Secs. 47; the present section is offered only as a back-of-the-envelope indication that those data sit in a density regime where partonic degrees of freedom are not phenomenologically unreasonable.
11 Why jet quenching is not visible — and why that is consistent
A legitimate question is: if a partonic medium is present in high-multiplicity $pp$, why is jet quenching not observed? Two independent, quantitative reasons make the absence of jet quenching compatible with, and indeed expected from, the present picture.
(i) Path-length scaling.
Radiative parton energy loss in a dense medium scales as
$\Delta E\propto\hat{q}\,L^{2},$ (4)
with $\hat{q}$ the transport coefficient and $L$ the path length. In central PbPb the medium extends over $L_{\mathrm{PbPb}}\sim 5$66 fm; in $pp$ the droplet extends over $L_{pp}\sim 1$22 fm. At comparable $\hat{q}$:
$\frac{\Delta E_{pp}}{\Delta E_{\mathrm{PbPb}}}\sim\left(\frac{L_{pp}}{L_{\mathrm{PbPb}}}\right)^{2}\sim\frac{1}{25}\text{--}\frac{1}{36}.$ (5)
An effect that is tens of GeV in PbPb becomes a few hundred MeV in $pp$. Current jet measurements in $pp$ are not sensitive to losses of that magnitude: the hadronic calorimetry threshold for identifying quenching is $\mathcal{O}(\text{GeV})$. The effect is therefore below the detection floor, not absent.
(ii) Quarkonium sensitivity scale.
A few hundred MeV are invisible to jet quenching but decisive for the $\Upsilon(3\mathrm{S})$, whose binding energy is $\sim 200$ MeV. A medium whose jet-quenching footprint is negligible is nonetheless capable of dissolving loosely bound quarkonia efficiently. The sensitivity of sequential $\Upsilon(n\mathrm{S})$ suppression to the density range
above the deconfinement crossover, below the jet-quenching detection threshold,
is structural. Had jet quenching been measured in $pp$ at $\varepsilon_{\mathrm{BJ}}\approx 1.7$ GeV/fm3, that would be in tension with our estimate and with the present interpretation. The experimental facts — strong sequential $\Upsilon$ suppression, strangeness enhancement, ridge, partonic $v_{2}$, no jet quenching — are all simultaneously consistent with a density regime in which $\Upsilon(n\mathrm{S})$ states, and not high-$p_{T}$ jets, are the uniquely sensitive probes.
12 Interpretation and closure
Four independent differential filters (cone, azimuth, sphericity, $p_{T}$-escape) and one temporal filter (non-prompt flat) act together on the data. Each filter places a different constraint on the admissible mechanisms. In their published forms, the hadronic and string-based frameworks considered in Sec. 8 fail at least one of these constraints without room for an obvious, parameter-free repair. On the evidence presented, the minimal set of structural features a surviving framework must carry is that of an early, globally correlated medium consistent with partonic degrees of freedom. The qualitative ingredients are:
  1. 1.
    Finite volume: core–corona with $R_{\mathrm{HBT}}\approx 1.8$ fm, a survival floor for corona quarkonia, and a central region that dissolves the excited states.
  2. 2.
    Momentum-dependent escape: crossing time $\tau_{\mathrm{cross}}\sim R/(\beta\gamma)$ becomes shorter than the medium lifetime at high $p_{T}$, generating the low-/high-$p_{T}$ asymmetry observed by LHCb.
  3. 3.
    Early action: the non-prompt flatness fixes the mechanism to act on the pre-resonance $b\bar{b}$ (or $c\bar{c}$) stage, compatible with Debye-like colour screening.
  4. 4.
    Equation-of-state consistency: the Bjorken-type estimate of the energy density (understood as a consistency check, not a measurement) sits in the soft-to-hard transition band identified by Campanini & Ferri [18], above the lattice-QCD crossover [20, 21].
The onset co-occurrence with strangeness enhancement and ridge, and the subsequent turning on of partonic $v_{2}$ at higher multiplicity, form a coherent multi-observable picture that is difficult to accommodate as a chance superposition of unrelated effects.
What this analysis does and does not claim.
This is a constraint paper. What it establishes, to the extent that the public CMS and LHCb datasets allow, is that the multi-differential structure of the measured $\Upsilon(n\mathrm{S})$ and $\psi(2\mathrm{S})/J/\psi$ patterns places strong joint requirements on any candidate mechanism, and that the hadronic and string-based frameworks we have been able to examine do not naturally satisfy those requirements simultaneously. The analysis should be read as a map of what any future framework must reproduce, and as an experimental case — not a formal demonstration — that the most economical reading of the combined evidence involves an early, globally correlated, partonic-like medium.
Data availability and source attribution
CMS 7 TeV data: HEPData ins1805867 (doi:10.17182/hepdata.95684.v1); LHCb 13 TeV $\psi(2\mathrm{S})/J/\psi$ data: JHEP 05 (2024) 243, Fig. 2 (CC BY 4.0, reproduced visually in Fig. 7). All numerical values in this paper come from the public datasets cited; no new measurement is claimed.